OpenStudy (anonymous):

5/3(3c-3)=1/5(20c+50) solve for c step by step please thanks

OpenStudy (mimi_x3):

lolss , hello againss , i will help yu agains xDD

OpenStudy (anonymous):

lol thanks sorry about ur question but its too hard and dont want to get it wrong

OpenStudy (mimi_x3):

what question ? i didnt post any questions lmao

OpenStudy (anonymous):

oh oops lol i thought you were mimi35 lmao my bad ;p

OpenStudy (mimi_x3):

LOLS , tht person stole my name :P

OpenStudy (anonymous):

lol ikr Dx

OpenStudy (mimi_x3):

5/(9c-9) = 1/(100c+250) 5(100c + 250) = 9c - 9 500c +1250 = 9c-9 491c = -1259 c = -1259/491 ( im not sure if im right though, substitite c into the equation to check if its right xDD)

OpenStudy (anonymous):

sorry i dont think ur right D;

OpenStudy (mimi_x3):

let me check it again

OpenStudy (mimi_x3):

ohh yeah my bad

OpenStudy (mimi_x3):

then what's the correct answer , im sure its right

OpenStudy (anonymous):

\[\frac{5}{3}(3x-3)=\frac{1}{5}(20x+50)\] \[5(x-1)=4x+10\] \[5x-5=4x+10\] \[x=15\]